OpenStudy (anonymous):

How do I find the twelfth term of 5,8,11,14,...

OpenStudy (anonymous):

add 3 until u get to the 12th term.

OpenStudy (anonymous):

ex- 14+3,17+3,20+3. until you get to the 12th term

OpenStudy (anonymous):

is there a formula that you should use

OpenStudy (anonymous):

well on that certain question, its +3 so just look for a pattern.

OpenStudy (anonymous):

Yeah: 1st term=5 2nd term=5+3 3rd term=5+2(3) 4th term=5+3(3) 12th term=5+11(3)=5+33=38